I have no idea how to approach this, can you please help me.
The question is as followed,
Flying to Kampala with a tailwind a plane averaged 185 km/h. On the return trip the plane only averaged 143 km/h while flying back into the same wind. What is the speed of the wind and the speed of the plane while in the air?
They are asking me to write a linear equation and solve.
X= could represent the speed of the wind and Y=the speed of the plane while in the air.
not sure how to get the answer


Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)


Flying to Kampala with a tailwind a plane averaged 185 km/h. On the return trip the plane only averaged 143 km/h while flying back into the same wind. What is the speed of the wind and the speed of the plane while in the air?
They are asking me to write a linear equation and solve.
X= could represent the speed of the wind and Y=the speed of the plane while in the air.
not sure how to get the answer
=======================
The plane's airspeed is the average of the 2 groundspeeds.
= (185+143)/2 = 164 km/hr
================
Windspeed is the difference between airspeed and groundspeed.
164 - 143 = 21 km/hr



Answer by ikleyn(52847)

.

It is a typical "tailwind and headwind" word problem.

See the lessons
    - Wind and Current problems
    - Wind and Current problems solvable by quadratic equations
    - Selected problems from the archive on a plane flying with and against the wind
in this site, where you will find other similar solved problems with detailed explanations.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the section "Word problems",  the topic "Travel and Distance problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.